LBC Verification and Checkout after Summer Shutdown 2013

Questions?? Please contact Michelle Edwards @ 520-661-0015 (anytime!)

Still to be done (9/10/2013):

Tilt Measurement

Filter Offsets

Guiding Check (Added 9/10/2013)


Summer 2013 Shutdown Work

Software

LBC version 1.4:
  • don't pop-up with "... system encountered a problem" if it is stopped because a STOP was requested (IT 3289)
  • don't disable filter motor when stopped - updates for new hardware
  • added Z4 to GUI (IT 4642)
  • logging rotator reduced trajectories
  • add setting DD entries for rotator angle and filter name (using testDouble and testString in TCS 2013B) (IT 4728)
  • reset guiding parameters when preset beginning, don't write guide.gif from IATrackFirst (IT 4411)
  • log an ERROR if background is zero on guide image (IT 3759, IT 4642)
  • misc mods to the test programs to make things more clear and some things easier

Hardware

  • Vacuum and cryo chek of both cryostat.
  • Mechanical upgrade of filter wheels with the new components using friction pads instead of electro brakes

Detector

  • LBC RED science camera fixed about vertical high frequency noise pattern at high flux levels. (IT 4754)
  • LBC BLU and RED cameras general check about ccd tech and science detector performances
  • Long tests for data acquisition taking more than 2000 frames for science and 8000 for tech images (IT 2631)



Summary of 2013 Post-Shutdown LBC Tasks

Daytime Checks (High Priority) [1.25 hr]:*DONE*

Basic start-up (0.75 hr) :

  • Check basic functionality and stability of cameras
  • Verify that the time on the CMU is correct and synchronized with our GPS * ssh telescope@lbccontrol * run "usr/sbin/ntpdate -q gps0" from the CMU. If the offset is less than a few milliseconds it is fine.
  • Check the load on CMU with top.

Check stability of cameras and data flow: (0.5 hr)

  • Obtain biases and a dark. OBs for this are in the Calib_OBs package: Calib_OBs/BIASDARK and are 25Bias_Bino.ob and 2Darks_Bino.ob.
  • Check the images as they appear in RB_Science. Note the median levels for each chip
  • Confirm that data are flowing to /newdata and the archive correctly.

Daytime Checks (Low Priority) [2 hr]

Check filter wheels (2 hr) TBD

  • Run subarray tests to find combination of dome lighting and exp times that illuminate every filter in closed dome situation
  • Create OBs that reflect these exposure times
  • This both checks the filter wheels thoroughly after a major mechanical changes AND will result in a streamlined way to check filters in daytime

Science check of software release v. 1.4 (0.1 hour) CHECKED -- although continue to try and note

  • Stop system purposefully several times during OBs to determine if pop-up no longer appears when STOP is requested (IT 3289)

Nighttime Checks (High Priority) [4.5]

Test dofpia, guiding, dithering (1 hr) DONE

  • Choose field with many stars to ensure guide star on tech chip. Useful targets: Calib_OBs/MLE_Test_targets/2013Aug30
  • Run dofpia and check outputs carefully
  • Check/correct pointing/co-pointing with lbcrangebal
  • Run OB with several dithers, filters, and exp times (Calib_OBs/MLE_Test_targets/2013Aug30). *Confirm offsets are sent to telescope and appear as expected on image
    • Confirm that guiding is working correctly

Test guiding after update of guiding executables (Added Sept 10th)

  • Choose field with many stars to ensure guide star on tech chip. Useful OBs: Calib_OBs/MLE_Test_targets/2013Aug30/Guide_Check/
  • Run dofpia and check outputs carefully
  • Check/correct pointing/co-pointing with lbcrangebal
  • Run script with 5 X 360s exposures to check guiding

Pointing check (1 hr) DONE

  • OBs in /home/LBTO/RB_VV_PointingOBs/
  • Slew to pointing star
  • Use LBTtools.Observe.lbcrangebal to correct pointing and co-pointing.
  • Go to more pointing stars and check pointing corrections

Measure rotator center (0.5 hr) DATA TAKEN IN ~1". TO BE ANALYZED; IF POSSIBLE ANOTHER SET IN BETTER CONDITIONS

  • Slew to a star field at moderate elevation and collimate. Useful targets: Calib_OBs/MLE_Test_targets/2013Aug30
  • Take a pair of 5-sec images at two diametrically opposed position angles (0 and 180, e.g.). Good idea to obtain at least two pairs.
    • Ask the operator to slew.
    • OBs are rotatorcenter_pa0.ob and rotatorcenter_pa180.ob (these do not slew the telescope) in /home/LBTO/Calib_OBs/ROTCEN_Check/
  • If there is time during the night, run script LBTtools.LBC.rotcenter to compute rotator centers.
    • Update eight header template files, blue_ehu_sci{1,2,3,4}.hdr and red_ehu_sci{1,2,3,4}.hdr on CMU: /home/lbccontrol/conf/ by updating the values for CRPIX1 and CRPIX2 in lines 2 and 3.
    • Define:
      • Delta_X = RotCen_X(new) - RotCen_X(old) and
      • Delta_Y = RotCen_Y(new) - RotCen_Y(old)
      • Then for chips 1,2,3:
        • CRPIX1(new) = CRPIX1(old) + Delta_X and
        • CRPIX2(new) = CRPIX2(old) + Delta_Y
      • And for chip 4:
        • CRPIX1(new) = CRPIX1(old) + Delta_Y and
        • CRPIX2(new) = CRPIX2(old) - Delta_X

Measure focal plane tilt using superfoc (1 hr)

  • Slew to a field with many stars, though not too dense. Useful targets: Calib_OBs/MLE_Test_targets/2013Aug30
  • Obtain focus sequences for at least 2 position angles (0 and 180, e.g.)
  • OBs are in/home/LBTO/Calib_OBs/SUPERFOC/:
    • RB_rVsuperfoc.ob7, RB_rVsuperfoc7_pa90.ob, RB_rV_superfoc7_pa180.ob and RB_rV_superfoc7_pa270.ob
    • These OBs step the primary mirror by 0.06 mm at 7 positions centered around the best-focus.

Measure filter focus offsets for g-sloan (0.5 hr)

* Procedure for obtaining these data:
    • Slew to a field with many stars, though not too dense. Useful targets: Calib_OBs/MLE_Test_targets/2013Aug30
    • Run dofpia
    • Run 3-4 superfoc sequences (/home/LBTO/Calib_OBs/SUPERFOC/) using any filter pair that includes g'
    • Run 3-4 superfoc sequencds that include the reference pair using the rV filters (RB_rV_superfoc.ob)*

Nighttime Checks: (Low Priority) [4-5 hr]:

Science check of software release v. 1.4 (0.5 hr)

  • confirm Z4 on GUI with any OB (IT 4642)
  • run an OB with a short exposure and U filter to ensure no guide star
    • check that guiding image reflects lack of guide star (IT 4411)
    • check for any errors in case of zero background ((IT 3759, IT 4642; likely hard to reproduce)

Flat fields (2-3 hr in twilight)

  • Obtain flats in as many pairs as possible of blue/red filters at two PAs. These are useful to:
    • check that there is nothing vignetting the field
    • build up a database of twilight sky flats for sciops and science use
  • Flat field OBs are created by mkskyflat1.pl in /home/LBTO/Calib_OBs/Mkskyflat1/

Measure filter focus offsets (2 hr)

* Procedure for obtaining these data:
    • Slew to a field with many stars, though not too dense. Useful targets: Calib_OBs/MLE_Test_targets/2013Aug30
    • Run dofpia
    • Run a few superfoc sequences (/home/LBTO/Calib_OBs/SUPERFOC/) using filter pairs
    • Run 3-4 superfoc sequencds that include the reference pair using the rV filters (RB_rV_superfoc.ob)*

-- MichelleEdwards - 27 Aug 2013
Topic revision: r7 - 10 Sep 2013, MichelleEdwards
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback